OpenStudy (anonymous):

If Anyone Helps I'll Become A Fan And Give Out A Medal Please Help! Solve. 4 t plus 6 is greater than or equal to ten Which graph shows the solution?

OpenStudy (anonymous):

Which graph shows the solution? A number line is numbered zero through ten. A solid circle is over the 1 and an arrow extends from this circle to the right. A number line is numbered zero through ten. A solid circle is over the 4 and an arrow extends from this circle to the right. A number line is numbered ten through twenty. An open circle is over the twelve and an arrow extends from this circle to the right. A number line is numbered ten through twenty. An open circle is over the sixteen and an arrow extends from this circle to the right.

OpenStudy (anonymous):

\[4 t + 6 \ge 10 \rightarrow 4 t \ge 10-6 \rightarrow 4 t \ge 4 \rightarrow t \ge 1\]

OpenStudy (anonymous):

A number line is numbered zero through ten. A solid circle is over the 1 and an arrow extends from this circle to the right.